OpenStudy (anonymous):

evaluated the integral if it converges. show diverges otherwise (poster in the threat )

OpenStudy (anonymous):

A. \[\int\limits_{\infty}^{-1} (e^(-3x)) dx\]

OpenStudy (anonymous):

B.\[\int\limits_{0}^{\infty} 3xe^{-4x^{2}} dx\]

OpenStudy (anonymous):

C. int_{0}^{infty} 4/ 1+9x^{2} dx

OpenStudy (anonymous):

\[\int\limits_{4}^{1} 4/{(x-1)^{2}}\]

OpenStudy (anonymous):

e. \[\int\limits_{4}^{1} x/{x^{2}-1} dx\]

OpenStudy (anonymous):

show the work

OpenStudy (anonymous):

only 5 question so i need to find whether those fiver are converges or diverges

OpenStudy (anonymous):